OpenStudy (anonymous):

Take 9 from 6,10 from 9,50 from 40, and leave 6. How come?????

OpenStudy (anonymous):

@AidenS think you can do this one

OpenStudy (anonymous):

9-6-? how do you do it?

OpenStudy (dan815):

check the letters

OpenStudy (dan815):

check roman numerals

OpenStudy (dan815):

check different bases

OpenStudy (thegamer007):

SIX - 9 (IX) = S 9 (IX) - 10 (X) = I 40 (XL) - 50 (L) = X => SIX

OpenStudy (traceur55):

(6-9)+(9-10))-(40-50)= (-3)+(-1)-(-10)= (-4)+10= 6 Or nah? Also, yes you can also do it with numerals. n It will spell out "SIX."